ANALYSIS OF THE FIRST HIGH RESOLUTION FTS INFRARED SPECTRA OF F_{2}^{11}BOH: THE ν_{8}, ν_{9}, 2ν_{9} AND ν_{4} BANDS.

\underline{A. PERRIN}, J.-M. FLAUD, Laboratoire de Photophysique Moléculaire, CNRS, Université Paris Sud, Campus d’Orsay, Bat 350, 91405 Orsay Cedex, France; H. BÜRGER, D.COLLET, Anorganische Chemie, FB 9, Universität-Gesamthochschule, D-42097 Wuppertal, Germany; J.DEMAISON, Laboratoire PhLAM, CNRS, Université de Lille I, Bat. P5, 59655 Villeneuve d’Ascq Cedex, France.

The reactive F_{2}BOH molecule was first detected by microwave spectroscopy ^{a}, and very recently observed by matrix IR spectroscopy ^{b}. We report here the first high resolution infrared study of F_{2}BOH. F_{2}BOH has been produced in a slow flow of (^{11}B and natural) BF_{3} through a glass tube filled of some specks of quartz (all using stainless stell equipment) on which water had been deposited. The IR spectrum has been recorded from 400 cm^{−1} to 1600 cm^{−1} at high resolution ( 2-3 x 10^{−3} cm^{−1}) using the Wuppertal Bruker 120 HR interferometer equipped with a cell of 1.2 m path length. In addition, ground state parameters have been determined from recent microwave measurements performed in Lille ^{c}. Among the recorded infrared bands, two c-type out-of-plane fundamental bands ν_{8} (BF_{2} bend) and ν_{9} (OH torsion) located at 684.16 cm^{−1} and 522.86 cm^{−1}, respectively, were analysed using a simple Watson -type Hamiltonian. The a/b type hybrid 2ν_{9} and ν_{4} bands centered at 1042.87 and 961.49 cm^{−1} were also studied. The analysis of ν_{4} (OH bending mode) was complicated by the existence of ”classical” vibrational rotational resonances linking the 4^{1} energy levels with those of the 7^{1}9^{1} dark overtone state. More surprising is the fact that both in the 2ν_{9} and ν_{4} bands, the P- and R-lines exhibit a regular doublets structure (of about 0.005 and 0.003 cm^{−1} respectively) which indicates the existence of large amplitude motions in the F_{2}BOH molecule.

aH. Takeo and R. F. Curl, J. Chem. Phys. 56, 4314 (1972)

bM. E. Jacox, K. K. Irikura , and W. E. Thompson, J. Chem. Phys.113, 5705 (2000)

cJ. Demaison, J.F.D’Eu et al. private communication